OpenStudy (anonymous):

If a fair coin is tossed three times, what is the probability of getting 3 tails ?

OpenStudy (lgbasallote):

TTT i believe it becomes \[\frac{1}{2} \times \frac{1}{2} \times \frac{1}{2}\]